OpenStudy (anonymous):

The ratio of girls to boys who participated in the quiz bowl was 7:5. There were 42 girls in the competition. How many boys participated? A) 28 B) 30 C) 35 D) 49

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

our proportion is: \[\frac{ 7 }{ 5 }=\frac{ 42 }{ x }\] where x is the number of boys.

OpenStudy (michele_laino):

So: \[7:5=42:x\] and applying the fundamental property of proportion, we have:

OpenStudy (anonymous):

so the answer is....

OpenStudy (anonymous):

lol

OpenStudy (michele_laino):

\[210=7*x\] x=30 so B.

OpenStudy (anonymous):

thankyou so much @Michele_Laino

OpenStudy (michele_laino):

thank you! @muneeb180